OpenStudy (anonymous):

Solve for x in the proportion x over the quantity x plus 12 equals 1 over the quantity x minus 1. x ≈ −6.02 and x ≈ 5.02 x ≈ −3.46 and x ≈ 3.46 x ≈ −3.14 and x ≈ 4.14 x ≈ −2.61 and x ≈ 4.61

OpenStudy (anonymous):

could u help me with this

OpenStudy (anonymous):

Solve for x in the proportion x/x+12=1/x-1

OpenStudy (lgbasallote):

\[\frac{x}{x+12} = \frac{1}{x-1}\] cross multiply \[x(x-1) = x + 12\] distribute \[x^2 - x = x = 12\] put everything on one side \[x^2 - x - x - 12 = 0\] \[x^2 - 2x - 12 = 0\] solve for x

OpenStudy (anonymous):

would it be a, b, c or d?